8 coins are tossed. What is the probability of getting at least one head showing on any of the coins?
Found 2 solutions by greenestamps, math_tutor2020:
Answer by greenestamps(13367)



P(at least one head) = 1 - P(no heads) = 1 - (1/2)^8 = 1 - 1/256 = 255/256

ANSWER: 255/256


Answer by math_tutor2020(3838)


We'll assume each coin is a fair coin. This means each coin has equal probability of getting heads or tails. That probability being 1/2.

Since each coin toss is independent of one another, the probability of getting all tails is (1/2)^8 = 1/256

The probability of at least one head showing up would then be 1 - (1/256) = 255/256
This works because we either get all tails, or at least one head shows up. The term "at least" means "that amount or more".

Answer: 255/256